OpenStudy (anonymous):

Solve the following system of equations using the substitution method x+3x=11 xy=8

OpenStudy (anonymous):

4x= 11 x =11/4 = 2.75 substitute x in xy= 8 2.75 * y =8 y =8/2.75 = 2.90

OpenStudy (anonymous):

i need an answer for x and y

OpenStudy (anonymous):

yes the above is answer for X and y x = 2.75 y= 2.90